OpenStudy (anonymous):

does cos(2/n) converge?

OpenStudy (zarkon):

\[\lim_{n\to\infty}\cos(2/n)=\cos(0)=1\]

OpenStudy (anonymous):

well, as n goes to infinity, or to zero?

OpenStudy (anonymous):

the sequence \[\frac{2}{n}\] converges to 0 as n gets larger and larger, so the sequence: \[\cos(\frac{2}{n}) \rightarrow \cos(0) = 1\]

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

i can never remember what the restrictions are for converging or diverging

OpenStudy (anonymous):

It diverges based off the nth term test for divergence. Which is what Zarkon wrote. The only way a series can converge is if the lim is zero. \[\sum_{n=1}^{\infty}a_n\] converges if and ONLY if \[\lim_{n \rightarrow \infty}a_n=0\]

OpenStudy (anonymous):

i think hes talking about the sequence, not the series.

OpenStudy (zarkon):

that is not true

OpenStudy (anonymous):

For series, it is true. It is not true for sequences.

OpenStudy (anonymous):

Well, that is misleading I suppose. The limit can be zero and it can still diverge. But I mean, if it converges, the lim is zero.

OpenStudy (zarkon):

\[\lim_{n\to\infty}a_n=0\] is a necessary condition not a sufficient condition

OpenStudy (anonymous):

That's what I meant.

OpenStudy (anonymous):

hokiefan, do you got to virginia tech?